OpenStudy (djbreezy):

Which is an equation in point-slope form for the given point and slope? Point: (7,4) ; Slope: 6 A. y-7=6(x+4) B. y-4=6(x-7) C. y-4=6(x+7) D. y+4=6(x-7)

OpenStudy (photon336):

This is all you need \[(y-y_{1}) = m(x-x_{1})\]

OpenStudy (photon336):

where m = slope. and y1 and x1 are your two points so based off of that which one would it be?

OpenStudy (djbreezy):

I think its A or C @Photon336

OpenStudy (texaschic101):

no... y - y1 = m(x - x1) m = 6 (7,4)....y1 = 4 and x1 = 7 now just sub...what do u get ?

OpenStudy (texaschic101):

y - y1...so whats y1 ?

OpenStudy (anonymous):

The answer is B y-4 = 6(x-7)

OpenStudy (texaschic101):

y1 = 4 so we have y - 4 y - 4 = m....so whats m ( the slope) ?

OpenStudy (djbreezy):

6

OpenStudy (texaschic101):

yes... so we have y - 4 = 6(x - x1)....so whats x1 ?

OpenStudy (texaschic101):

x1 = 7 answer is y- 4 = 6(x - 7) any questions ?

OpenStudy (djbreezy):

no

OpenStudy (djbreezy):

thx

OpenStudy (texaschic101):

yw